theyadoreshayyy:

A cylindrical piece of iron pipe is shown below. The wall of the pipe is 1.25 inch thick: What is the approximate inside volume of the pipe? 88 cubic inches 49 cubic inches 142 cubic inches 154 cubic inches

theyadoreshayyy:

Ladidie:

whats this assignment called maybe i can go on quizzez and find the awnser

ItzAgar:

Which ones do you think we can eliminate?

ItzAgar:

@theyadoreshayyy you there?

ItzAgar:

Oh she offline :/.

Florisalreadytaken:

there are 2 walls right? the left, and the right side one. thats why we multiply 1.25 inch by 2: \[ 2 \times 1.25 = 2.5 in \] lets find the diameter of the base of the smaller circle \( 6-2.5=3.5 \) that means the radi of the smaller circle is \[ \frac{3.5}{2}\Rightarrow r=1.75 \] now lets just find the volume of it just like we would normaly do: \[ V= \pi r^2 \times h \] \[ V= \pi \times 1.75^2 \times 16 \] \[ V= \pi \times 3.0625 \times 16 \] \[ V= \pi \times 49 \] \[ V\approx 154 \ in^3 \]